Please help me solve:
Susan Maricano invested part of her $18,000 bonus in a fund that paid an 8% profit and invested the rest in stock that suffered a 3% loss. Find the amount of each investment if her over all net profit was $780.
The amount invested at 8% is:
The amount invested in stock is:
This is what I have: her net profit is 780 total amount investing is 18,000 profit of 8% and loss of 3%. x=8% and y=-3.
I tried to use this formula: .08x-0.3y=780 but then i got totally confused and stuck. Please help.

Answer by checkley77(12844)

.08X-.03(18,000-X)=780
.08X-540+.03X=780
.11X=780+540
.11X=1,320
X=1,320/.11
X=$12,000 AMOUNT INVESTED @ 8%.
18,000-12,000=$6,000 INVESTED @ 3%
PROOF:
.08*12,000-.03*6,000=780
960-180=780
780=780